OpenStudy (anonymous):

solve for x: log 7x = log 56

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

base is 7 or 10?

myininaya (myininaya):

nvm i think its 10

OpenStudy (anonymous):

if the base is not specified it is 10

myininaya (myininaya):

log7+logx=log56 logx=log56-log7 logx=log(56/7) logx=log(8) x=8

OpenStudy (anonymous):

\[7x=56\] \[x=8\]

myininaya (myininaya):

yes im a retard lol

OpenStudy (anonymous):

doesn't matter what the base is, only that logs are one to one yes?

OpenStudy (anonymous):

thats ok myinaya you should have seen me square both sides yesterday

myininaya (myininaya):

i know im a retard

OpenStudy (anonymous):

when i could have just factored

myininaya (myininaya):

lol

OpenStudy (anonymous):

LOL...thanks to both of you

OpenStudy (anonymous):

actually it looks like you are an answering machine!

myininaya (myininaya):

lol i am but aimstre is beating me badly

OpenStudy (anonymous):

yeah well i will never catch up.

OpenStudy (anonymous):

how does he do it?

myininaya (myininaya):

he hacks

OpenStudy (anonymous):

for one thing he (she?) is fast. for another he is here 24/7

myininaya (myininaya):

jk

OpenStudy (anonymous):

multiple accounts?